# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v1

@@ Appendix B Hyperbolic Kernel Derivatives
By stacking the row derivatives $\frac{\partial}{\partial\bm{x}}\bm{K}_{i}\in\mathbb{R}^{1\times 4\times 4}$ along the first dimension and the column derivatives $\frac{\partial}{\partial\bm{z}}\bm{K}_{j}\in\mathbb{R}^{4\times 1\times 4}$ along the second, we get the full third-order kernel derivatives $\frac{\partial}{\partial\bm{x}}\bm{K},\frac{\partial}{\partial\bm{z}}\bm{K}\in\mathbb{R}^{4\times 4\times 4}$. As discussed in Section 4.2, automatic differentiation tools such as PyTorch cannot evaluate the helper functions $g(u)$, $h(u)$, and $q(u)$ as $u$ approaches $-1$ from below. To address this issue, we replace the general function definitions with their analytical limits, when $u$ gets too close to $-1$. This is equivalent to the hyperbolic distance between the two kernel inputs falling below a predefined threshold value $\rho=\text{arccosh}(-u)=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})\leq 1\mathrm{e}{-4}$. The analytical limit expressions for the helper functions are given as
$\lim_{u\rightarrow-1^{-}}g(u)$ $=\lim_{u\rightarrow-1^{-}}\left(\frac{2\rho^{2}}{\kappa s^{2}}-\frac{1}{s^{2}}-\frac{u\rho}{s^{3}}\right)$ $=\frac{2}{\kappa}+\dfrac{1}{3}\,,$ (59)
$\lim_{u\rightarrow-1^{-}}h(u)$ $=\lim_{u\rightarrow-1^{-}}\left(\dfrac{\text{d}}{\text{d}u}g(u)+\frac{g(u)2\rho}{\kappa s}\right)$ $=\dfrac{4}{\kappa^{2}}+\dfrac{6}{3\kappa}+\dfrac{4}{15}\,,$ (60)
$\lim_{u\rightarrow-1^{-}}q(u)$ $=\lim_{u\rightarrow-1^{-}}\left(\dfrac{\text{d}}{\text{d}u}h(u)+\frac{h(u)2\rho}{\kappa s}\right)$ $=\dfrac{8}{\kappa^{3}}+\dfrac{8}{\kappa^{2}}+\dfrac{14}{5\kappa}+\dfrac{12}{35}\,.$ (61)
Appendix C Experimental Details and Additional Results
@@
Apps. C.1-C.4 provide additional details regarding the data and GPHLVM training process for each of the experiments described in Section 5. App C.5 provides additional insights into the hand grasps experiment of Section 5.4.
C.1 C-shape
@@
The proof-of-concept experiment on the $\mathsf{C}$-shape dataset was designed as a minimalist example with a simple setup. The training data consists of $N=1000$ two-dimensional points arranged in a $\mathsf{C}$ shape and scaled to fit within the unit circle. In the Euclidean case, these datapoints are used as both the latent points and the observations, so that $\bm{X}=\bm{Y}\in\mathbb{R}^{1000\times 2}$. In the hyperbolic case, the datapoints can be interpreted as elements of the Poincaré model. We represent each latent point and observation in the Lorentz model using the Poincaré to Lorentz mapping $\bm{x}_{\mathcal{L}}=\frac{1}{1-\|\bm{x}\|^{2}}\begin{bmatrix}1+\|\bm{x}\|^{2}&2\bm{x}^{\mathsf{T}}\end{bmatrix}^{\mathsf{T}}\in\mathbb{H}^{2}_{\mathcal{L}}$. In this proof-of-concept experiment, we manually set the kernel variance, lengthscale, and noise variance of both GPLVM and GPHLVM to $(\tau,\kappa,\sigma^{2}_{y})=(0.7,0.15,0.69)$. Additionally, for the GPHLVM, we used the 2D hyperbolic heat kernel with $L=3000$ rejection samples. We used constant zero mean for both models.
@@
The start- and end-points of the geodesics shown in Figure 1 correspond to one of the $\mathsf{C}$-shape latent points. The geodesics are represented by $M=25$ regularly-spaced points. Geodesics on the base Euclidean manifold correspond to straight lines from $\bm{x}_{0}$ to $\bm{x}_{M-1}$. In the hyperbolic case, the geodesics on the base manifold are given as $\bm{x}_{i}=\text{Exp}_{\bm{x}_{0}}(t_{i}\text{Log}_{\bm{x}_{0}}(\bm{x}_{M-1}))$ with $t_{i}=i/(M-1)$ and $i=\{0,...,M-1\}$. The geodesics on the base manifolds were used as initialization for the pullback geodesic optimization. To compute the pullback geodesic, we applied Riemannian Adam for $200$ steps with a learning rate of $0.005$, optimizing the combined loss function (22) based on both curve energy and spline energy, as introduced in Section 3.3. In this experiment, the curve energy and spline energy were weighted equally, i.e., $\lambda=1$ in Equation (21).
@@
Finally, we generated a $110\times 110$ grid of points to visualize the volume of the pullback metric, filtering out those points lying outside the unit circle in the hyperbolic case. Then, we evaluated the pullback metric tensor at each remaining grid point. In the Euclidean case, we computed the volume at each grid point $\bm{x}^{*}$ by evaluating the determinant of the $2\times 2$ pullback metric tensor, $\sqrt{\text{det}(\bm{G}_{\bm{x}^{*}})}$. In the hyperbolic case, the pullback metric tensor is a $3\times 3$ matrix that lies on the $2$-dimensional tangent space, leading to one of its eigenvalues being zero. To address this, we compute the volume as the product of the two nonzero eigenvalues.
C.2 MNIST Digits
@@
C.3 Multi-cellular Robot Design
We consider multi-cellular robots composed of a $5\times 5$ grid of cells. Possible cell types are empty, solid, soft, a horizontal actuator or a vertical actuator. We generate a dataset of $216$ multicellular robots following the hierarchical approach of Dong et al., (2024), which we explain next for completeness. The root nodes of the hierarchy are two coarse robots composed of only vertical and only horizontal actuator cells. A component is defined as a group of cells of the same type, i.e., the root robots have a single component. Children nodes of each robot are designed by dividing large components into $2$ smaller components using K-means and by assigning a different type of cell to one of these components. This process is repeated iteratively to obtain robots with up to $8$ components. This results in a tree of multi-cellular robots, with coarse-grained robots close to the roots and fine-grained robots at the leaves.
To account for the hierarchical structure associated with the tree structure of the multi-cellular robots, we incorporate graph-based priors in our GPHLVM as proposed by Jaquier et al., (2024). This is achieved by leveraging the stress loss,
@@ C.3 Multi-cellular Robot Design
C.4 Hand Grasps Generation
For the hand grasp experiment, we used data from the KIT Whole-Body Motion Database (Mandery et al.,, 2016). The dataset consists of $38$ trajectories, with motion recordings from two human subjects (IDs $2122$ and $2123$). Each subject performed $19$ different grasp types [^2] The three-finger sphere grasp of the GRASP taxonomy (Feix et al.,, 2016) is missing from the dataset. of the GRASP taxonomy (Feix et al.,, 2016). Each grasp consists of the subject grasping an object from a table, lifting it, and placing it back.
@@ C.4 Hand Grasps Generation
In contrast to the previous experiments, this experiment involves trajectory data. To preserve the trajectory structure during training, we incorporated a dynamics prior $p(\bm{X}_{2:N}\mid\bm{X}_{1:N-1})$, similar to Wang et al., (2008), but using the wrapped Gaussian distribution (36). We also used back constraints $\bm{X}=k(\bm{Y},\bm{Y})\bm{C}$, which allow for a smooth mapping from the observation space to the latent space. The back-constraints kernel is defined as a Euclidean SE kernel with variance $\tau=1$ and a lengthscale $\kappa=0.4$. Moreover, each grasp type is identified with a lead node of the quantitative taxonomy of hand grasps (Stival et al.,, 2019). More specifically, the first point of each trajectory, which corresponds to the resting state, is assigned to the root node $c_{0}$ of the taxonomy graph, and the last point to the taxonomy node representing the corresponding grasp type. The taxonomy node $c_{14}$, for instance, represents the stick grasp. The number of edges between different grasp types in the taxonomy graph defines a distance function, $\text{d}_{\mathbb{G}}(c_{i},c_{j})$. To account for the hierarchical structure associated with the taxonomy, we additionally incorporate graph-based priors in our GPHLVM as proposed by Jaquier et al., (2024). This is achieved by leveraging the stress loss (62) to match the pairwise latent distances with the graph distances. During training, we minimize the stress loss over the first and last latent points of each trajectory along with maximizing the GPHLVM marginal likelihood by optimizing,
@@
Figure 6: Time-series plots of all $24$ dimensions of the joint space introduced in Figure 4. The left and right columns show the decoded geodesics of the 2D and 3D latent spaces. The top and bottom rows show the mean of the decoded hyperbolic and pullback geodesics ( and ) and their uncertainty as a gray envelope. For reference, the same training trajectory to the spherical grasp ( ) and reversed training trajectory from the ring grasp ( ) are included.
C.5 Additional Hand Grasps Results
@@
This section provides further insights into the hand grasps experiment described in Section 5.4. First, we present additional visualizations of the decoded geodesics. While Figure 4 only showed $3$ dimensions of the joint space, Figure 6 depicts the decoded geodesics across all $24$ dimensions, i.e., over the $24$ degrees of freedom (DoF) of the hand model introduced in Mandery et al., (2016). This model counts two joints for the wrist, four joints for each finger, and two additional joints for the ring and little fingers to allow a better fist closure.
@@
Second, we discuss the influence of initialization on optimizing the pullback geodesics. As described in Section C.4, we initialize the pullback geodesic with two concatenated hyperbolic geodesics: One from the ring grasp to the origin and another from the origin to the spherical grasp. This initialization incorporates knowledge about the expected path in the taxonomy graph. In contrast, Figure 7 shows the optimized pullback geodesics when the optimization is initialized solely with a hyperbolic geodesic from the ring grasp to the spherical grasp. In the two-dimensional case, we observe that the optimized pullback geodesic path differs significantly from the one shown in Figure 4. This discrepancy is likely due to the geodesic optimization getting stuck in a local minimum, which may be a consequence of the Monte Carlo approximation of the 2D hyperbolic heat kernel. Interestingly, the three-dimensional pullback geodesic does not exhibit this issue, further supporting the hypothesis that the Monte Carlo approximation may be responsible since the 3D case does not rely on this approximation. In the 3D case, the optimized pullback geodesic follows a path very similar to the one shown in Figure 4.
@@
Figure 7: Hand grasps transition when initializing the geodesic optimization with the hyperbolic geodesic. As in Figure 4 the optimized geodesic on the pullback metric ( ) and the hyperbolic geodesic ( ), transition from a ring grasp ( ) to a spherical grasp ( ). Top: 2D and 3D latent spaces of the trained models. Bottom: Time-series plots of $3$ dimensions of the joint space for the decoded hyperbolic geodesic ( ) and pullback geodesic ( ).

### v2

@@ Appendix B Hyperbolic Kernel Derivatives
By stacking the row derivatives $\frac{\partial}{\partial\bm{x}}\bm{K}_{i}\in\mathbb{R}^{1\times 4\times 4}$ along the first dimension and the column derivatives $\frac{\partial}{\partial\bm{z}}\bm{K}_{j}\in\mathbb{R}^{4\times 1\times 4}$ along the second, we get the full third-order kernel derivatives $\frac{\partial}{\partial\bm{x}}\bm{K},\frac{\partial}{\partial\bm{z}}\bm{K}\in\mathbb{R}^{4\times 4\times 4}$. As discussed in Sec. 4.2, automatic differentiation tools such as PyTorch cannot evaluate the helper functions $g(u)$, $h(u)$, and $q(u)$ as $u$ approaches $-1$ from below. To address this issue, we replace the general function definitions with their analytical limits, when $u$ gets too close to $-1$. This is equivalent to the hyperbolic distance between the two kernel inputs falling below a predefined threshold value $\rho=\text{arccosh}(-u)=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})\leq 1\mathrm{e}{-4}$. The analytical limit expressions for the helper functions are given as
$\lim_{u\rightarrow-1^{-}}g(u)$ $=\lim_{u\rightarrow-1^{-}}\left(\frac{2\rho^{2}}{\nu s^{2}}-\frac{1}{s^{2}}-\frac{u\rho}{s^{3}}\right)$ $=\frac{2}{\nu}+\dfrac{1}{3}\,,$ (64)
$\lim_{u\rightarrow-1^{-}}h(u)$ $=\lim_{u\rightarrow-1^{-}}\left(\dfrac{\text{d}}{\text{d}u}g(u)+\frac{g(u)2\rho}{\nu s}\right)$ $=\dfrac{4}{\nu^{2}}+\dfrac{6}{3\nu}+\dfrac{4}{15}\,,$ (65)
$\lim_{u\rightarrow-1^{-}}q(u)$ $=\lim_{u\rightarrow-1^{-}}\left(\dfrac{\text{d}}{\text{d}u}h(u)+\frac{h(u)2\rho}{\nu s}\right)$ $=\dfrac{8}{\nu^{3}}+\dfrac{8}{\nu^{2}}+\dfrac{14}{5\nu}+\dfrac{12}{35}\,.$ (66)
Appendix C Hyperbolic Pullback Geodesics
In Sec. 3.3, we described the computation of geodesics according to the hyperbolic pullback metric introduced in Sec. 3.13.2. Algorithm 1 summarizes the computation of a hyperbolic pullback geodesic by minimizing (24) with Riemannian optimization.
Algorithm 1 Computation of hyperbolic pullback geodesics
  Input: Start and end points $\bm{x}_{1}$, $\bm{x}_{M}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$ of the pullback geodesic, pullback metric $\bm{G}^{\text{P},\mathcal{L}}$.
  Output: Pullback geodesic represented as a set of points $\{\bm{x}_{i}\}_{i=1}^{M}$ with $\bm{x}_{i}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$.
  Initialization:
  Initialize the pullback geodesic, e.g., as the hyperbolic geodesic from $\bm{x}_{1}$ to $\bm{x}_{M}$.
  Discretize the pullback geodesic with $M$ points $\{\bm{x}_{i}\}_{i=1}^{M}$.
  Geodesic computation:
  repeat
     Compute the curve energy $E$ (23).
     Compute the spline energy $E_{\text{spline}}$.
     ${\bm{x}_{2},...,\bm{x}_{M-1}}\leftarrow\mathsf{RiemannianOptStep}(E+\lambda E_{\text{spline}})$.
  until convergence
Appendix D Experimental Details and Additional Results
@@
Apps. D.1-D.4 provide additional details regarding the data and GPHLVM training process for each of the experiments described in Sec. 5. App D.5 provides additional insights into the hand grasps experiment of Sec. 5.4.
D.1 C-shape
@@
The proof-of-concept experiment on the $\mathsf{C}$-shape dataset was designed as a minimalist example with a simple setup. The training data consists of $N=1000$ two-dimensional points arranged in a $\mathsf{C}$ shape and scaled to fit within the unit circle. In the Euclidean case, these datapoints are used as both the latent points and the observations, so that $\bm{X}=\bm{Y}\in\mathbb{R}^{1000\times 2}$. In the hyperbolic case, the datapoints can be interpreted as elements of the Poincaré model. We represent each latent point and observation in the Lorentz model using the Poincaré to Lorentz mapping $\bm{x}_{\mathcal{L}}=\frac{1}{1-\|\bm{x}\|^{2}}\begin{bmatrix}1+\|\bm{x}\|^{2}&2\bm{x}^{\mathsf{T}}\end{bmatrix}^{\mathsf{T}}\in\mathbb{H}^{2}_{\mathcal{L}}$. In this proof-of-concept experiment, we manually set the kernel variance, lengthscale, and noise variance of both GPLVM and GPHLVM to $(\tau,\kappa,\sigma^{2}_{y})=(0.7,0.15,0.69)$. Additionally, for the GPHLVM, we used the 2D hyperbolic SE kernel with $L=3000$ rejection samples. We used constant zero mean for both models.
@@
The start- and end-points of the geodesics shown in Fig. 2 correspond to one of the $\mathsf{C}$-shape latent points. The geodesics are represented by $M=25$ regularly-spaced points. Geodesics on the base Euclidean manifold correspond to straight lines from $\bm{x}_{0}$ to $\bm{x}_{M-1}$. In the hyperbolic case, the geodesics on the base manifold are given as $\bm{x}_{i}=\text{Exp}_{\bm{x}_{0}}(t_{i}\text{Log}_{\bm{x}_{0}}(\bm{x}_{M-1}))$ with $t_{i}=i/(M-1)$ and $i=\{0,...,M-1\}$. The geodesics on the base manifolds were used as initialization for the pullback geodesic optimization. To compute the pullback geodesic, we applied Riemannian Adam for $200$ steps with a learning rate of $0.005$, optimizing the combined loss function (24) based on both curve energy and spline energy, as introduced in Sec. 3.3. In this experiment, the curve energy and spline energy were weighted equally, i.e., $\lambda=1$ in Eq. (23).
@@
Finally, we generated a $110\times 110$ grid of points to visualize the volume of the pullback metric, filtering out those points lying outside the unit circle in the hyperbolic case. Then, we evaluated the pullback metric tensor at each remaining grid point. In the Euclidean case, we computed the volume at each grid point $\bm{x}^{*}$ by evaluating the determinant of the $2\times 2$ pullback metric tensor, $\sqrt{\text{det}(\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}^{*}})}$. In the hyperbolic case, the pullback metric tensor is a $3\times 3$ matrix that lies on the $2$-dimensional tangent space, leading to one of its eigenvalues being zero. To address this, we compute the volume as the product of the two nonzero eigenvalues.
D.2 MNIST Digits
@@
D.3 Multi-cellular Robot Design
We consider multi-cellular robots composed of a $5\times 5$ grid of cells. Possible cell types are empty, solid, soft, a horizontal actuator or a vertical actuator. We generate a dataset of $216$ multicellular robots following the hierarchical approach of Dong et al. (2024), which we explain next for completeness. The root nodes of the hierarchy are two coarse robots composed of only vertical and only horizontal actuator cells. A component is defined as a group of cells of the same type, i.e., the root robots have a single component. Children nodes of each robot are designed by dividing large components into $2$ smaller components using K-means and by assigning a different type of cell to one of these components. This process is repeated iteratively to obtain robots with up to $8$ components. This results in a tree of multi-cellular robots, with coarse-grained robots close to the roots and fine-grained robots at the leaves.
To account for the hierarchical structure associated with the tree structure of the multi-cellular robots, we incorporate graph-based priors in our GPHLVM as proposed by Jaquier et al. (2024). This is achieved by leveraging the stress loss,
@@ D.3 Multi-cellular Robot Design
D.4 Hand Grasps Generation
For the hand grasp experiment, we used data from the KIT Whole-Body Motion Database (Mandery et al., 2016). The dataset consists of $38$ trajectories, with motion recordings from two human subjects (IDs $2122$ and $2123$). Each subject performed $19$ different grasp types [^1] The three-finger sphere grasp of the GRASP taxonomy (Feix et al., 2016) is missing from the dataset. of the GRASP taxonomy (Feix et al., 2016). Each grasp consists of the subject grasping an object from a table, lifting it, and placing it back.
@@ D.4 Hand Grasps Generation
In contrast to the previous experiments, this experiment involves trajectory data. To preserve the trajectory structure during training, we incorporated a dynamics prior $p(\bm{X}_{2:N}\mid\bm{X}_{1:N-1})$, similar to Wang et al. (2008), but using the wrapped Gaussian distribution (38). We also used back constraints $\bm{X}=k(\bm{Y},\bm{Y})\bm{C}$, which allow for a smooth mapping from the observation space to the latent space. The back-constraints kernel is defined as a Euclidean SE kernel with variance $\tau=1$ and a lengthscale $\kappa=0.4$. Moreover, each grasp type is identified with a lead node of the quantitative taxonomy of hand grasps (Stival et al., 2019). More specifically, the first point of each trajectory, which corresponds to the resting state, is assigned to the root node $c_{0}$ of the taxonomy graph, and the last point to the taxonomy node representing the corresponding grasp type. The taxonomy node $c_{14}$, for instance, represents the stick grasp. The number of edges between different grasp types in the taxonomy graph defines a distance function, $\text{d}_{\mathbb{G}}(c_{i},c_{j})$. To account for the hierarchical structure associated with the taxonomy, we additionally incorporate graph-based priors in our GPHLVM as proposed by Jaquier et al. (2024). This is achieved by leveraging the stress loss (67) to match the pairwise latent distances with the graph distances. During training, we minimize the stress loss over the first and last latent points of each trajectory along with maximizing the GPHLVM marginal likelihood by optimizing,
@@
Figure 7: Time-series plots of all $24$ dimensions of the joint space introduced in Fig. 5. The left and right columns show the decoded geodesics of the 2D and 3D latent spaces. The top and bottom rows show the mean of the decoded hyperbolic ( ) and pullback geodesics ( ) and their uncertainty as a gray envelope. For reference, the same training trajectory to the spherical grasp ( ) and reversed training trajectory from the ring grasp ( ) are included.
(a) GPLVM, $\mathbb{R}^{2}$
(b) GPLVM, $\mathbb{R}^{3}$
Figure 8: Hand grasps transition from a ring ( ) to a spherical grasp ( ) using Euclidean ( ) and pullback ( ) geodesics in (a) $\mathbb{R}^{2}$ and (b) $\mathbb{R}^{3}$. Top left: Latent embeddings. The latent trajectory points are colored according to their corresponding grasp class. The background color in $\mathbb{R}^{2}$ represents the pullback metric volume. Top right: Time-series plots of $2$ dimensions of the joint space. The top and bottom rows show the mean of the decoded hyperbolic ( ) and pullback geodesics ( ) with their uncertainty as a gray envelope. A training trajectory to the spherical grasp ( ) and a reversed training trajectory from the ring grasp ( ) are included for reference. Bottom: Generated hand trajectories from the decoded hyperbolic (top) and pullback (bottom) geodesics.
D.5 Additional Hand Grasps Results
@@
This section provides further insights into the hand grasps experiment described in Sec. 5.4. First, we present additional visualizations of the decoded geodesics. While Fig. 5 only showed $2$ dimensions of the joint space, Fig. 7 depicts the decoded geodesics across all $24$ dimensions, i.e., over the $24$ degrees of freedom (DoF) of the hand model introduced in Mandery et al. (2016). This model counts two joints for the wrist, four joints for each finger, and two additional joints for the ring and little fingers to allow a better fist closure.
Second, we compare the GPHLVM and geodesics generated in Sec. 5.4 to their counterpart Euclidean models. We train two GPLVMs augmented with a taxonomy graph prior and a dynamic prior similarly as the GPHLVMs of Sec. 5.4. Fig. 8 show the resulting 2D and 3D latent spaces, alongside the Euclidean and Euclidean pullback geodesic interpolating from a ring to a spherical grasps and their corresponding decoded hand motions.
@@
Third, we discuss the influence of initialization on optimizing the pullback geodesics. As described in Sec. D.4, we initialize the pullback geodesic with two concatenated hyperbolic geodesics: One from the ring grasp to the origin and another from the origin to the spherical grasp. This initialization incorporates knowledge about the expected path in the taxonomy graph. In contrast, Fig. 9 shows the optimized pullback geodesics when the optimization is initialized solely with a hyperbolic geodesic from the ring grasp to the spherical grasp. In the two-dimensional case, we observe that the optimized pullback geodesic path differs significantly from the one shown in Fig. 5. This discrepancy is likely due to the geodesic optimization getting stuck in a local minimum, which may be a consequence of the Monte Carlo approximation of the 2D hyperbolic SE kernel. Interestingly, the three-dimensional pullback geodesic does not exhibit this issue, further supporting the hypothesis that the Monte Carlo approximation may be responsible since the 3D case does not rely on this approximation. In the 3D case, the optimized pullback geodesic follows a path very similar to the one shown in Fig. 5.
@@
Figure 9: Hand grasps transition when initializing the geodesic optimization with the hyperbolic geodesic. As in Fig. 5, a transition from a ring ( ) to a spherical grasp ( ) is obtained using hyperbolic ( ) and pullback ( ) geodesics in $\mathbb{H}^{2}_{\mathcal{L}}$ (left) and $\mathbb{H}^{3}_{\mathcal{L}}$ (right). Top: 2D and 3D latent spaces of the trained models. Bottom: Time-series plots of $3$ dimensions of the joint space for the decoded hyperbolic geodesic ( ) and pullback geodesic ( ).